5. You grow a culture of E. coli to an OD reading of 0.8. Your instructor wants 40 colonies on a plate. Ideally, you want to only spread 100ul on a LB plate. How would you carry out the proper dilutions in order to arrive at 40 cells on a plate? Please show your calculations and all work. (1 O.D. = 1 x 108 cells/mL).
Answer #1

Volume plated = 100uL

Number of colonies =40 colonies

Step 1: Convert O.D to cfu/ml

(O.D of 0.8) x (1x 10^8) / (1 O.D) = 8 x10^ 7 cfu/ml

Step 2: Number of bacteria/mL= Number of colonies * Dilution of the sample

  Number of colonies * Volume plated * Total dilutions used

8 x10^ 7 cfu/ml = 40 colonies * 0.1 ml * Total dilutions

Total dilutions=   8 x10^ 7 cfu/ml) / ( 40 colonies) * (0.1 ml) = 32 X10^7